# Hybrid Feedback for Three-dimensional Convex Obstacle Avoidance (Extended version)

### v2

@@ V Stability Analysis
$\underbrace{\begin{matrix}\mathbf{\dot{x}}\\ \mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\end{matrix}\begin{matrix}[l]\;=\mathbf{u}(\xi)\\ \;=\mathbf{0}\\ \;=\mathbf{0}\\ \;=0\\ \;=1\end{matrix}}_{\dot{\xi}=\mathbf{F}(\xi),{\xi\in\mathcal{F}}}\quad\underbrace{\begin{matrix}\mathbf{x}^{+}\\ \begin{bmatrix}\mathbf{h}^{+}\\ \mathbf{a}^{+}\\ m^{+}\\ s^{+}\end{bmatrix}\end{matrix}\begin{matrix}[l]=\mathbf{x}\vspace{0.56cm}\\ \vspace{0.7cm}\in\mathbf{L}(\xi)\end{matrix}}_{\xi^{+}\in\mathbf{J}(\xi),{\xi\in\mathcal{J}}},$ (32)
@@
The next lemma shows that the hybrid closed-loop system (32) satisfies the hybrid basic conditions [16, Assumption 6.5], which guarantees the well-posedness of the hybrid closed-loop system.
@@ VI Implementation Procedure
We consider a workspace with convex obstacles that satisfies Assumption 8 with some $\bar{r}_{s}>0$, as discussed in Section III. The target location is set at the origin within the interior of the obstacle-free workspace $\mathcal{W}_{r_{a}}^{\circ}$. The center of the robot is initialized in the obstacle-free workspace $\mathcal{W}_{r_{a}}$. The variables $\mathbf{h},\mathbf{a},m$ and $s$ are initialized in the sets $\mathcal{W}_{r_{a}},\mathbb{S}^{2},\mathbb{M}$ and $\mathbb{R}_{\geq 0}$, respectively. The parameter $\gamma\in(0,\bar{r}_{s}-r_{s})$ enables the algorithm to avoid one obstacle at a time, as discussed in Remark 3. The parameters $\gamma_{a}$ and $\gamma_{s}$ are set to satisfy $0<\gamma_{a}<\gamma_{s}<\gamma.$ A sufficiently small value for $\bar{\epsilon}$ is selected as stated in Lemma 1, and the parameter $\epsilon\in(0,\bar{\epsilon}].$ We set a sensing radius $R_{s}>r_{a}+\gamma$ such that the robot can detect the boundary of the obstacles within its line-of-sight inside the region $\mathcal{B}_{R_{s}}(\mathbf{x})$.
@@
2:  Initialize $\mathbf{x}(0,0)\in\mathcal{W}_{r_{a}}$, $\mathbf{h}(0,0)\in\mathcal{W}_{r_{a}}$, $\mathbf{a}(0,0)\in\mathbb{S}^{2}$, $m(0,0)\in\mathbb{M}$ and $s(0,0)\in\mathbb{R}_{\geq 0}$. Choose sufficiently small value of $\bar{\epsilon}$ according to Lemma 1, and initialize $\epsilon\in(0,\bar{\epsilon}].$ Identify the parameter $\bar{r}_{s}$, as discussed in Section III, and set parameters $\gamma,\gamma_{s}$ and $\gamma_{a}$ such that $0<\gamma_{a}<\gamma_{s}<\gamma<\bar{r}_{s}-r_{s}$. Choose $R_{s}>r_{a}+\gamma$, used in (34).
@@ VII Simulation Results
In the first simulation, we consider a single three-dimensional convex obstacle, as shown in Fig. 2. The initial location of the robot $\mathbf{x}(0,0)$ is represented by the green diamond symbol, and the origin is selected as the target location. The radius of the robot $r=0.15m$ and the safety distance $r_{s}=0.1m$. The parameter $\gamma=0.4m$ and the parameter $\epsilon$, used in (23), is set to $0.1m$. The gains $\kappa_{s}$ and $\kappa_{r}$, used in (12a), are set to $1$ and $0.5$, respectively. The sensing radius $R_{s}$, used in (34), is set to $1m$. The simulations are performed in MATLAB 2020a. In Fig. 2, the red-colored trajectory represents the motion of the robot in the move-to-target mode, whereas the blue-colored trajectory represents the motion of the robot in the obstacle-avoidance mode. When the robot switches from the move-to-target mode to the obstacle-avoidance mode, the algorithm updates the coordinates of the hit point $\mathbf{h}$ and the unit vector $\mathbf{a}$ according to (28). Then, the robot moves in the $\gamma-$neighborhood of the nearest obstacle with its center in the hyperplane $\mathcal{P}(\mathbf{h},\mathbf{a})$, as shown in Fig. 2. When the center of the robot enters in the jump set of the obstacle-avoidance mode, it switches back to the move-to-target mode and converges to the target location at the origin.
@@
We consider an unbounded workspace i.e., obstacle $\mathcal{O}_{0}=\emptyset$, with $2$ three-dimensional, convex obstacles, as shown in Fig. 3. We apply the proposed hybrid feedback controller (12) for robot initialized at $8$ different locations in the obstacle-free workspace. The target is located at the origin. The radius of the robot is set to $0.15m$. The minimum safety distance $r_{s}=0.05m$ and the parameter $\gamma=0.2m$. The gains $\kappa_{s}$ and $\kappa_{r}$, used in (12a), are set to $0.5$ and $1$, respectively. The sensing radius $R_{s}$, used in (34), is set to $1m$. The parameter $\epsilon$, used in (23), is set to 0.05m. From Fig. 3, it can be observed that the robot converges to the target location while simultaneously avoiding collisions with the obstacles. Fig. 4 shows that the center of the robot stays at least $r_{a}$ meters away from the boundary of the obstacles. The complete simulation video can be found at https://youtu.be/67eDVXH1wbw.
@@ IX-C Proof of Lemma 3
First we prove that the union of the flow and jump sets covers exactly the obstacle-free state space $\mathcal{K}$. For $m=0$, according to (19) and (20), by construction we have $\mathcal{F}_{0}^{\mathcal{W}}\cup\mathcal{J}_{0}^{\mathcal{W}}=\mathcal{W}_{r_{a}}.$ Similarly, for $m=1$, according to (22) and (24), by construction one has $\mathcal{F}_{m}^{\mathcal{W}}\cup\mathcal{J}_{m}^{\mathcal{W}}=\mathcal{W}_{r_{a}}$. Inspired by [12, Appendix 11], the satisfaction of the following equation:
@@
Now, inspired by [12, Appendix 1], for the hybrid closed-loop system (32), with data $\mathcal{H}=(\mathcal{F},\mathbf{F},\mathcal{J},\mathbf{J}),$ define $\mathbf{S}_{\mathcal{H}}(\mathcal{K})$ as the set of all maximal solutions $\xi$ to $\mathcal{H}$ with $\xi(0,0)\in\mathcal{K}$. Since $\mathcal{F}\cup\mathcal{J}=\mathcal{K}$, each $\xi\in\mathbf{S}_{\mathcal{H}}(\mathcal{K})$ has range rge $\text{rge }\xi\subset\mathcal{K}.$ Additionally, if every maximal solution $\xi\in\mathbf{S}_{\mathcal{H}}(\mathcal{K})$ is complete, then the set $\mathcal{K}$ will be forward invariant [18, Definition 3.13]. Since the hybrid closed-loop system (32) satisfies the hybrid basic conditions, as stated in Lemma 2, one can use [16, Proposition 6.10], to verify the following viability condition:
@@
Hence, according to [16, Proposition 6.10], since (37) holds for all $\xi\in\mathcal{F}\setminus\mathcal{J}$, there exists a nontrivial solution to $\mathcal{H}$ for each initial condition in $\mathcal{K}$. Finite escape time can only occur through flow. They can neither occur for $\mathbf{x}$ in the set $\mathcal{F}_{1}^{\mathcal{W}}$, as this set is bounded as per definition (24), nor for $\mathbf{x}$ in the set $\mathcal{F}_{0}^{\mathcal{W}}$ as this would make $\mathbf{x}^{\intercal}\mathbf{x}$ grow unbounded, and would contradict the fact that $\frac{d}{dt}(\mathbf{x}^{\intercal}\mathbf{x})\leq 0$ in view of the definition of $\mathbf{u}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)$. Therefore, all maximal solutions do not have finite escape times. Furthermore, according to (32), $\mathbf{x}^{+}=\mathbf{x},$ and from the definition of the update law in (28) and (31), it follows immediately that $\mathbf{J}(\mathcal{J})\subset\mathcal{K}$. Hence, the solutions to the hybrid closed-loop system (32) cannot leave $\mathcal{K}$ through jump and, as per [16, Proposition 6.10], all maximal solutions are complete.
@@ IX-D Proof of Theorem 1
Forward invariance and stability: The forward invariance of the obstacle-free set $\mathcal{K}$, for the hybrid closed-loop system (32), is immediate from Lemma 3. We next prove the stability of $\mathcal{A}$ using [16, Definition 7.1].
Since $\mathbf{0}\in(\mathcal{W}_{r_{a}})^{\circ}$, there exists $\mu_{1}>0$ such that $\mathcal{B}_{\mu_{1}}(\mathbf{0})\cap(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))^{\circ}=\emptyset.$ According to (19), there exists $\mu_{2}>0$ such that $\mathcal{B}_{\mu_{2}}(\mathbf{0})\cap\mathcal{J}_{0}^{\mathcal{W}}=\emptyset$. We define the set $\mathcal{S}_{\mu}:=\{\xi\in\mathcal{K}|\mathbf{x}\in\mathcal{B}_{\mu}(\mathbf{0})\},$ where $\mu\in(0,\min\{\mu_{1},\mu_{2}\}).$ Notice that for all initial conditions $\xi(0,0)\in\mathcal{S}_{\mu}$, the control input, after at most one jump corresponds to the move-to-target mode and it steers the state $\mathbf{x}$ towards the origin according to the control input vector $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x},\kappa_{s}>0$. Hence, for each $\mu\in(0,\min\{\mu_{1},\mu_{2}\}),$ the set $\mathcal{S}_{\mu}$ is forward invariant for the hybrid closed-loop system (32).
Consequently, for every $\rho>0$, one can choose $\sigma\in(0,\min\{\mu_{1},\mu_{2},\rho\})$ such that for all initial conditions $\xi(0,0)$ with $d(\xi(0,0),\mathcal{A})\leq\sigma$, one has $d(\xi(t,j),\mathcal{A})\leq\rho$ for all $(t,j)\in\text{ dom }\xi,$ where $d(\xi,\mathcal{A})^{2}=\underset{(\mathbf{0},\bar{\mathbf{h}},\bar{\mathbf{a}},\bar{m},s)\in\mathcal{A}}{\inf}(\lVert\mathbf{x}\rVert^{2}+\lVert\mathbf{h}-\bar{\mathbf{h}}\rVert^{2}+\lVert\mathbf{a}-\bar{\mathbf{a}}\rVert^{2}+(m-\bar{m})^{2}+(s-\bar{s})^{2})=\lVert\mathbf{x}\rVert^{2}.$ Hence, according to [18, Definition 3.1], the target set $\mathcal{A}$ is stable for the hybrid closed-loop system (32). Next, we proceed to establish the convergence properties of the set $\mathcal{A}$.
Attractivity: We aim to show that for the proposed hybrid closed-loop system (32), the target set $\mathcal{A}$ is globally attractive in the set $\mathcal{K}$ using [18, Defintion 3.1 and Remark 3.5]. In other words, we prove that for all initial conditions $\xi(0,0)\in\overline{\mathcal{F}}\cup\mathcal{J}=\mathcal{K}$, every maximal solution $\xi$ to the hybrid closed-loop system is complete and satisfies
@@
According to Lemma 4, there exists $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1)$ with $t_{2}<\infty$ such that $\xi(t_{2},j_{1}+1)\in\mathcal{J}_{1}$. Notice that, according to (22) and (23), one has $\|\mathbf{x}(t_{2},j_{1}+1)\|<\|\mathbf{x}(t_{1},j_{1}+1)\|$. In other words, according to Lemma 4, the proposed navigation algorithm ensures that, at the instance where the the control switches from the obstacle-avoidance mode to the move-to-target mode, the origin is closer to the point $\mathbf{x}$ than to the last point where the control switched to the obstacle-avoidance mode. Furthermore, when the control input corresponds to the move-to-target mode, it steers the state $\mathbf{x}$ towards the origin under the influence of control $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x},\kappa_{s}>0.$ Consequently, given that the workspace $\mathcal{W}$ and the obstacles $\mathcal{O}_{i},i\in\mathbb{I}\setminus\{0\}$, are compact, it can be concluded that the solution $\xi(t,j)$ will contain finite number of jumps and will satisfy (44).
@@
Hence, the target set $\mathcal{A}$ is globally attractive in the set $\mathcal{K}$ for the proposed hybrid closed-loop system (32). In addition, since the set $\mathcal{A}$ is stable for the hybrid closed-loop system (32), it is globally asymptotically stable in the set $\mathcal{K}$ for the hybrid closed-loop system (32) as per [18, Remark 3.5].

### v3

@@ V STABILITY ANALYSIS
$\underbrace{\begin{matrix}\mathbf{\dot{x}}\\ \mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\end{matrix}\begin{matrix}[l]\;=\mathbf{u}(\xi)\\ \;=\mathbf{0}\\ \;=\mathbf{0}\\ \;=0\\ \;=1\end{matrix}}_{\dot{\xi}=\mathbf{F}(\xi),\;{\xi\in\mathcal{F}},}\quad\underbrace{\begin{matrix}\mathbf{x}^{+}\\ \begin{bmatrix}\mathbf{h}^{+}\\ \mathbf{a}^{+}\\ m^{+}\\ s^{+}\end{bmatrix}\end{matrix}\begin{matrix}[l]=\mathbf{x}\vspace{0.56cm}\\ \vspace{0.7cm}\in\mathbf{L}(\xi)\end{matrix}}_{\xi^{+}\in\mathbf{J}(\xi),\;{\xi\in\mathcal{J}},}$ (31)
@@
The next lemma shows that the hybrid closed-loop system (31) satisfies the hybrid basic conditions [23, Assumption 6.5], which guarantees the well-posedness of the hybrid closed-loop system.
@@ VI IMPLEMENTATION PROCEDURE
We consider a workspace with convex obstacles that satisfies Assumption 8 with some $\delta>0$. The target location is set at the origin within the interior of the obstacle-free workspace $\mathcal{W}_{r_{a}}^{\circ}$. The center of the robot is initialized in the obstacle-free workspace $\mathcal{W}_{r_{a}}$. The variables $\mathbf{h},\mathbf{a},m$ and $s$ are initialized in the sets $\mathcal{W}_{r_{a}},\mathbb{S}^{2},\mathbb{M}$ and $\mathbb{R}_{\geq 0}$, respectively. Since $\gamma\in(0,\delta-r_{s})$, the algorithm to avoid one obstacle at a time, as discussed in Remark 3. The parameters $\gamma_{a}$ and $\gamma_{s}$ are set to satisfy $0<\gamma_{a}<\gamma_{s}<\gamma.$ A sufficiently small value for $\bar{\epsilon}$ is selected as stated in Lemma 1, and the parameter $\epsilon\in(0,\bar{\epsilon}].$ We set a sensing radius $R_{s}>r_{a}+\gamma$ such that the robot can detect the boundary of the obstacles within its line-of-sight inside the region $\mathcal{B}_{R_{s}}(\mathbf{x})$.
@@
2:  Initialize $\mathbf{x}(0,0)\in\mathcal{W}_{r_{a}}$, $\mathbf{h}(0,0)\in\mathcal{W}_{r_{a}}$, $\mathbf{a}(0,0)\in\mathbb{S}^{2}$, $m(0,0)\in\mathbb{M}$ and $s(0,0)\in\mathbb{R}_{\geq 0}$. Choose sufficiently small value of $\bar{\epsilon}$ according to Lemma 1, and initialize $\epsilon\in(0,\bar{\epsilon}].$ Identify the parameter $\delta$, as discussed in Assumption 8, and set parameters $\gamma,\gamma_{s}$ and $\gamma_{a}$ such that $0<\gamma_{a}<\gamma_{s}<\gamma<\delta-r_{s}$. Choose $R_{s}>r_{a}+\gamma$, used in (33). Set $\tau_{s}>0$ and $\delta_{s}\in(0,\tau_{s})$.
@@ VII SIMULATION RESULTS
In the first simulation, we consider a single three-dimensional convex obstacle, as shown in Fig. 2. The initial location of the robot $\mathbf{x}(0,0)$ is represented by the green diamond symbol, and the origin is selected as the target location. The radius of the robot $r=0.15m$ and the safety distance $r_{s}=0.1m$. The parameter $\gamma=0.4m$ and the parameter $\epsilon$, used in (22), is set to $0.1m$. The gains $\kappa_{s}$ and $\kappa_{r}$, used in (11a), are set to $1$ and $0.5$, respectively. The sensing radius $R_{s}$, used in (33), is set to $2m$. The simulations are performed in MATLAB 2020a. In Fig. 2, the red-colored trajectory represents the motion of the robot in the move-to-target mode, whereas the blue-colored trajectory represents the motion of the robot in the obstacle-avoidance mode. When the robot switches from the move-to-target mode to the obstacle-avoidance mode, the algorithm updates the coordinates of the hit point $\mathbf{h}$ and the unit vector $\mathbf{a}$ according to (27). Then, the robot moves in the $\gamma-$neighborhood of the nearest obstacle with its center in the hyperplane $\mathcal{P}(\mathbf{h},\mathbf{a})$, as shown in Fig. 2. When the center of the robot enters in the jump set of the obstacle-avoidance mode, it switches back to the move-to-target mode and converges to the target location at the origin.
@@
We consider an unbounded workspace i.e., obstacle $\mathcal{O}_{0}=\emptyset$, with $2$ three-dimensional, convex obstacles, as shown in Fig. 3. We apply the proposed hybrid feedback controller (11) for robot initialized at $8$ different locations in the obstacle-free workspace. The target is located at the origin. The radius of the robot is set to $0.15m$. The minimum safety distance $r_{s}=0.05m$ and the parameter $\gamma=0.2m$. The gains $\kappa_{s}$ and $\kappa_{r}$, used in (11a), are set to $0.5$ and $1$, respectively. The sensing radius $R_{s}$, used in (33), is set to $1m$. The parameter $\epsilon$, used in (22), is set to 0.05m. In Fig. 3, the dotted portion of the trajectory represents the motion of the robot in the move-to-target mode, whereas the solid portion corresponds to the movement in the obstacle-avoidance mode. From Fig. 3, it can be observed that the robot converges to the target location while simultaneously avoiding collisions with the obstacles. Fig. 4 shows that the center of the robot stays at least $r_{a}$ meters away from the boundary of the obstacles.
@@
We compare our approach with the separating hyperplane method developed in [13] which can be implemented in a priori unknown environments using information from a range-bearing sensor mounted on the robot. However, unlike our approach, it only works for strongly convex obstacles [13, Assumption 2]. The workspace contains two obstacles, an ellipsoidal obstacle (strongly convex) and a polygonal obstacle that does not satisfy [13, Assumption 2]. The robot’s radius is set to $0.2m$, with a minimum safety distance $r_{s}=0.1m$ and $\gamma=0.5m$. The gains $\kappa_{s}=0.4$ and $\kappa_{r}=2$ are used in (11a), while the sensing radius $R_{s}=2m$ is used in (33), and $\epsilon=0.5m$ in (22).
In Fig. 5a, the dotted portion of the trajectory represents the motion of the robot in the move-to-target mode, whereas the solid portion corresponds to the motion in the obstacle-avoidance mode. The proposed hybrid feedback controller guarantees safe, global asymptotic convergence to the target location at the origin starting from eight different initial locations as shown in Fig. 5a. In contrast, the separating hyperplane approach generates trajectories that converge to an undesired equilibrium (local minimum), as seen in Fig. 5b. The complete simulation video is available at222[Online]. Available: https://youtu.be/xihWpec8kis.
Figure 5: (a) Trajectories obtained using our proposed approach. (b) Trajectories obtained using the separating hyperplane approach [13].
@@ IX-C Proof of Lemma 3
First we prove that the union of the flow and jump sets covers exactly the obstacle-free state space $\mathcal{K}$. For $m=0$, according to (18) and (19), by construction we have $\mathcal{F}_{0}^{\mathcal{W}}\cup\mathcal{J}_{0}^{\mathcal{W}}=\mathcal{W}_{r_{a}}.$ Similarly, for $m=1$, according to (21) and (23), by construction one has $\mathcal{F}_{m}^{\mathcal{W}}\cup\mathcal{J}_{m}^{\mathcal{W}}=\mathcal{W}_{r_{a}}$. Inspired by [20, Appendix 11], the satisfaction of the following equation:
@@
Now, inspired by [20, Appendix 1], for the hybrid closed-loop system (31), with data $\mathcal{H}=(\mathcal{F},\mathbf{F},\mathcal{J},\mathbf{J}),$ define $\mathbf{S}_{\mathcal{H}}(\mathcal{K})$ as the set of all maximal solutions $\xi$ to $\mathcal{H}$ with $\xi(0,0)\in\mathcal{K}$. Since $\mathcal{F}\cup\mathcal{J}=\mathcal{K}$, each $\xi\in\mathbf{S}_{\mathcal{H}}(\mathcal{K})$ has range rge $\text{rge }\xi\subset\mathcal{K}.$ Additionally, if every maximal solution $\xi\in\mathbf{S}_{\mathcal{H}}(\mathcal{K})$ is complete, then the set $\mathcal{K}$ will be forward invariant [25, Definition 3.13]. Since the hybrid closed-loop system (31) satisfies the hybrid basic conditions, as stated in Lemma 2, one can use [23, Proposition 6.10], to verify the following viability condition:
@@
Hence, according to [23, Proposition 6.10], since (36) holds for all $\xi\in\mathcal{F}\setminus\mathcal{J}$, there exists a nontrivial solution to $\mathcal{H}$ for each initial condition in $\mathcal{K}$. Finite escape time can only occur through flow. They can neither occur for $\mathbf{x}$ in the set $\mathcal{F}_{1}^{\mathcal{W}}$, as this set is bounded as per definition (23), nor for $\mathbf{x}$ in the set $\mathcal{F}_{0}^{\mathcal{W}}$ as this would make $\mathbf{x}^{\intercal}\mathbf{x}$ grow unbounded, and would contradict the fact that $\frac{d}{dt}(\mathbf{x}^{\intercal}\mathbf{x})\leq 0$ in view of the definition of $\mathbf{u}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)$. Therefore, all maximal solutions do not have finite escape times. Furthermore, according to (31), $\mathbf{x}^{+}=\mathbf{x},$ and from the definition of the update law in (27) and (30), it follows immediately that $\mathbf{J}(\mathcal{J})\subset\mathcal{K}$. Hence, the solutions to the hybrid closed-loop system (31) cannot leave $\mathcal{K}$ through jump and, as per [23, Proposition 6.10], all maximal solutions are complete.
@@ IX-D Proof of Theorem 1
Forward invariance and stability: The forward invariance of the obstacle-free set $\mathcal{K}$, for the hybrid closed-loop system (31), is immediate from Lemma 3. We next prove the stability of $\mathcal{A}$ using [23, Definition 7.1].
Since $\mathbf{0}\in(\mathcal{W}_{r_{a}})^{\circ}$, there exists $\mu_{1}>0$ such that $\mathcal{B}_{\mu_{1}}(\mathbf{0})\cap(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))^{\circ}=\emptyset.$ According to (18), there exists $\mu_{2}>0$ such that $\mathcal{B}_{\mu_{2}}(\mathbf{0})\cap\mathcal{J}_{0}^{\mathcal{W}}=\emptyset$. Additionally, as per (21), there exists $\mu_{3}>0$ such that $\mathcal{B}_{\mu_{3}}(\mathbf{0})\subset\mathcal{J}_{1}^{\mathcal{W}}$. We define the set $\mathcal{S}_{\mu}:=\{\xi\in\mathcal{K}|\mathbf{x}\in\mathcal{B}_{\mu}(\mathbf{0})\},$ where $\mu\in(0,\min\{\mu_{1},\mu_{2},\mu_{3}\}).$ Notice that for all initial conditions $\xi(0,0)\in\mathcal{S}_{\mu}$, the control input, after at most one jump corresponds to the move-to-target mode and it steers the state $\mathbf{x}$ towards the origin according to the control input vector $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x},\kappa_{s}>0$. Hence, for each $\mu\in(0,\min\{\mu_{1},\mu_{2},\mu_{3}\}),$ the set $\mathcal{S}_{\mu}$ is forward invariant for the hybrid closed-loop system (31).
Consequently, for every $\rho>0$, one can choose $\sigma\in(0,\min\{\mu_{1},\mu_{2},\mu_{3},\rho\})$ such that for all initial conditions $\xi(0,0)$ with $d(\xi(0,0),\mathcal{A})\leq\sigma$, one has $d(\xi(t,j),\mathcal{A})\leq\rho$ for all $(t,j)\in\text{ dom }\xi,$ where $d(\xi,\mathcal{A})^{2}=\underset{(\mathbf{0},\bar{\mathbf{h}},\bar{\mathbf{a}},\bar{m},s)\in\mathcal{A}}{\inf}(\lVert\mathbf{x}\rVert^{2}+\lVert\mathbf{h}-\bar{\mathbf{h}}\rVert^{2}+\lVert\mathbf{a}-\bar{\mathbf{a}}\rVert^{2}+(m-\bar{m})^{2}+(s-\bar{s})^{2})=\lVert\mathbf{x}\rVert^{2}.$ Hence, according to [25, Definition 3.1], the target set $\mathcal{A}$ is stable for the hybrid closed-loop system (31). Next, we proceed to establish the convergence properties of the set $\mathcal{A}$.
Attractivity: We aim to show that for the proposed hybrid closed-loop system (31), the target set $\mathcal{A}$ is globally attractive in the set $\mathcal{K}$ using [25, Defintion 3.1 and Remark 3.5]. In other words, we prove that for all initial conditions $\xi(0,0)\in\overline{\mathcal{F}}\cup\mathcal{J}=\mathcal{K}$, every maximal solution $\xi$ to the hybrid closed-loop system is complete and satisfies
@@
According to Lemma 4, there exists $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1)$ with $t_{2}<\infty$ such that $\xi(t_{2},j_{1}+1)\in\mathcal{J}_{1}$. Notice that, according to (21) and (22), one has $\|\mathbf{x}(t_{2},j_{1}+1)\|<\|\mathbf{x}(t_{1},j_{1}+1)\|$. In other words, according to Lemma 4, the proposed navigation algorithm ensures that, at the instance where the control switches from the obstacle-avoidance mode to the move-to-target mode, the origin is closer to the point $\mathbf{x}$ than to the last point where the control switched to the obstacle-avoidance mode. Furthermore, when the control input corresponds to the move-to-target mode, it steers the state $\mathbf{x}$ towards the origin under the influence of control $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x},\kappa_{s}>0.$ Consequently, given that the workspace $\mathcal{W}$ and the obstacles $\mathcal{O}_{i},i\in\mathbb{I}\setminus\{0\}$, are compact, it can be concluded that the solution $\xi(t,j)$ will contain finite number of jumps and will satisfy (43).
@@
Hence, the target set $\mathcal{A}$ is globally attractive in the set $\mathcal{K}$ for the proposed hybrid closed-loop system (31). In addition, since the set $\mathcal{A}$ is stable for the hybrid closed-loop system (31), it is globally asymptotically stable in the set $\mathcal{K}$ for the hybrid closed-loop system (31) as per [25, Remark 3.5].